how long will a boy sitting near the window of a train travelling at54 km/hr see atrain passing by in the direction opposite to it with a speed of 36km/hr . the lenght of the slow moving train is 100m
Answer by lwsshak3(11628)

how long will a boy sitting near the window of a train travelling at54 km/hr see atrain passing by in the direction opposite to it with a speed of 36km/hr . the lenght of the slow moving train is 100m
**
let x= time the boy will see the passing train
100m=sum of the distances both trains travel while passing each other
54+36=90 km/hr is the effective speed traveling thru 100 m while passing each other
travel time
=distance/speed
=100/(90*10^3) hr
=(100/(90*10^3)*3600
= 4 seconds
time the boy will see the passing train=4 seconds